# Partial Semigroupoid Actions on Sets

### v1

@@ 4. Application in semigroups and categories
${{}_{s}Y}=(\varphi\circ\varphi^{-1})({{}_{s}Y})\subseteq\varphi({{}_{s}X})\subseteq{{}_{s}Y}\quad\therefore\quad\varphi({{}_{s}X})={{}_{s}Y},$
@@
Consequently, for each $s\in\mathcal{S}$, we have ${{}_{s}E}=E\setminus\delta(X_{0})$. On the other hand, since $(i,\gamma,\mathcal{S}^{1}\otimes X)$ is a globalization of $(\alpha,X)$, there is a unique morphism of partial actions $\varphi\colon(\beta,E)\to(\gamma,\mathcal{S}^{1}\otimes X)$ such that $\varphi\circ\delta=i$. This morphism is defined by
@@
Therefore, $\varphi\colon E\to\mathcal{S}^{1}\otimes X$ is a bijection. However, the inverse function $\varphi^{-1}\colon\mathcal{S}^{1}\otimes X\to E$ may not be a morphism of partial actions, since in $(\gamma,\mathcal{S}^{1}\otimes X)$ every element acts on $i(X_{0})=\varphi(\delta(X_{0}))$ as the identity function, while in $(\beta,E)$ no element can act on $\delta(X_{0})$.
We end this section noticing that the fact that any degenerate partial semigroup action admits a non-degenerate globalization that provides an initial object in the category $((\alpha,X)\to I^{\ast})$ is a convenient property of semigroups. Indeed, if $(\beta,Y)$ is a non-degenerate global action of $\mathcal{S}$, then ${{}_{s}Y}=Y$ for every $s\in\mathcal{S}$. Therefore, in order to define a morphism $\Phi\colon(\beta,E)\to(\theta,Y)$, there is no need to verify that $\Phi({{}_{s}E})\subseteq{{}_{s}Y}$. The next example illustrates, in the context of semigroupoids, how a degenerate partial action may fail to have such a non-degenerate globalization.
@@
${{}_{e}Y}=\{0,1,2\},\quad\beta_{e}(0,1,2)=(0,1,0),\quad{{}_{f}Y}=\{0,1,3\}\quad\text{and}\quad\beta_{f}(0,1,3)=(0,1,0).$
@@
In this case $(\alpha,X)$ is a degenerate global action of $S$. We know that $i\colon X\to Y$, given by $i(1,2)=(1,2)$, is an embedding. Notice that $j\colon X\to Y$, $j(1,2)=(1,3)$ is also an embedding.
Suppose that $(\delta,\theta,E)$ is a non-degenerate globalization of $(\alpha,X)$. Since $S^{s}\neq\emptyset$ for every $s\in S$, we obtain from Lemma 2.10(ii) that $E_{s}\subseteq{{}_{t}E}$, for some $t\in S^{s}$. Therefore, $E\subseteq\cup_{s\in S}{{}_{s}E}$. Now, the embedding $\delta\colon X\to E$ must satisfy $\delta(2)\in{{}_{e}E}$ or $\delta(2)\in{{}_{f}E}$. Suppose further that $(\delta,\theta,E)$ is initial object in the category $((\alpha,X)\to I^{\ast})$. Then there must be two morphisms of partial actions, $i^{\prime},j^{\prime}\colon(\theta,E)\to(\beta,Y)$, satisfying $i^{\prime}\circ\delta=i$ and $j^{\prime}\circ\delta=j$. Therefore
$i^{\prime}(\delta(2))=i(2)=2\in{{}_{e}Y}\setminus{{}_{f}Y}\quad\text{and}\quad j^{\prime}(\delta(2))=j(2)=3\in{{}_{f}Y}\setminus{{}_{e}Y}.$
Since $i^{\prime}$ is a morphism of partial actions, if $\delta(2)\in{{}_{f}E}$ then $i^{\prime}(\delta(2))\in{{}_{f}Y}$, a contradiction. So it must be $\delta(2)\in{{}_{e}E}$. But in this case, as $j^{\prime}$ is a morphism of partial actions, it must be $j^{\prime}(\delta(2))\in{{}_{e}Y}$, a contradiction. Therefore, the degenerate partial action $(\alpha,X)$ does not admit a non-degenerate globalization that provides an initial object for $((\alpha,X)\to I^{\ast})$.
@@
In this subsection, we investigate the relation between partial category actions and partial semigroupoid actions of a category and show that our construction of universal globalization coincides with the one presented in [21, §3].
Throughout this subsection, $\mathcal{C}$ will denote a category (recall our conventions in Definition 2.3). The following characterization of partial category actions was given in [21, Proposition 8].
@@
A partial action of $\mathcal{C}$ on a set $X$ is a pair ($(\{{{}_{g}X}\}_{g\in\mathcal{C}},\{\alpha_{g}\}_{g\in\mathcal{C}})$ consisting of a collections $\{{{}_{g}X}\}_{g\in\mathcal{C}}$ of subsets of $X$ and a collection $\{\alpha_{g}\}_{g\in\mathcal{C}}$ of functions $\alpha_{g}\colon{{}_{g}X}\to X$ such that:
@@
    $X=\cup_{e\in\mathcal{C}_{0}}X_{e}$ and $\alpha_{e}=id_{{}_{e}X}$, for all $e\in\mathcal{C}_{0}$;
@@
Compared to the original characterization, we omit the collection of subsets $\{X_{g}\}_{g\in\mathcal{C}}$, since it is implicitly defined as the images of the collection of functions $\{\alpha_{g}\}_{g\in\mathcal{C}}$, and interchange the roles of $X_{g}$ and ${{}_{g}X}$. This change aims the comparison with partial group actions, [11, Definition 1.2], where $\alpha_{g}\colon X_{g^{-1}}\to X_{g}$, or in our notation, $X_{g}=X_{g}$ and ${{}_{g}X}=X_{g^{-1}}$. It is straightforward from (c1) that any partial category action is non-degenerate.
Similarly to the relation between semigroups and monoids ([17, Lemma 2.4]), there is a particular class of partial semigroupoid actions that correspond to partial category actions. To describe it, we need the notion of an identity element in a semigroupoid.
@@
Notice that $S_{0}$ may be the empty set. Furthermore, for each $s\in S$, there is at most one identity $e\in S^{s}$. For if $e,f\in S_{0}\cap S^{s}$, then $s=se=(sf)e$. Therefore, $fe$ is defined and, hence, $e=f$.
@@
Let us start with (ii). Suppose that $(e,s)\in S^{(2)}$, then $es=s$. From condition (P1) we have $\alpha_{s}^{-1}({{}_{e}X}\cap X_{s})={{}_{s}X}\cap{{}_{es}X}={{}_{s}X}$. Applying $\alpha_{s}$ to this equality, we obtain ${{}_{e}X}\cap X_{s}=X_{s}$. Therefore $X_{s}\subseteq{{}_{e}X}$. Let $x\in{{}_{s}X}={{}_{s}X}\cap{{}_{es}X}$. Then condition (P2) implies $(\alpha_{e}\circ\alpha_{s})(x)=\alpha_{es}(x)=\alpha_{s}(x)$. That is, for every $y=\alpha_{s}(x)\in X_{s}$ we have $\alpha_{e}(y)=y\in X_{e}$. So $X_{s}\subseteq{{}_{e}X}\cap X_{e}$.
@@
Let $S$ be a semigroupoid and $(\alpha,X)$ be a partial action (Definition 2.9). We say that $(\alpha,X)$ is a categorical partial action of $S$ if the following conditions hold:
@@
Assume (i). To prove (ii) it remains to prove (C1). From the second part of (c1) we have $\alpha_{e}=id_{{}_{e}X}$, for every $e\in\mathcal{C}_{0}$. That is, ${{}_{e}X}=X_{e}$. From the first part of (c1) we obtain
@@
Suppose that (G2) holds, that is, ${{}_{h}X}={{}_{gh}X}$, whenever the composition $gh$ is defined. In particular, for $h=D(g)$, the composition $gD(g)$ exists and, hence, ${{}_{D(g)}X}={{}_{gD(g)}X}={{}_{g}X}$. So condition (c4) holds. Assume (c4) and notice that, for $g\in\mathcal{C}$,
@@
Therefore, if $\mathcal{C}^{g}=\mathcal{C}^{h}$, then $D(g)=D(h)$. In this case ${{}_{g}X}={{}_{D(g)}X}={{}_{h}X}$. So (G1) holds. ∎
@@
The notion of globalization presented in [21, Definition 11] is more general than ours. The term globalization means any morphism of partial actions $\delta\colon(\alpha,X)\to(\beta,Y)$ where $(\beta,Y)$ is a global action, while we require $\delta$ to be an embedding. On the other hand, universal globalization means a globalization $(\delta,\beta,Y)$ such that, for every morphism $\varphi\colon(\alpha,X)\to(\theta,Z)$ where $(\theta,Z)$ is a global action, there is a unique morphism $\Phi\colon(\beta,Y)\to(\theta,Z)$ satisfying $\Phi\circ\delta=\phi$. Apart from the definition of globalization, the idea of universal globalization agree with ours.
@@
    In [21, Theorem 4] it has been proved that $(\alpha,X)$ admits a universal globalization $(i,\gamma,Y)$. Furthermore, [21, Remark 22] shows that $i$ is, in fact, an embedding.
@@
    From (1) and Proposition 2.23(ii), it follows that, for any universal globalization $(\delta,\beta,Z)$ of $(\alpha,X)$, $\delta$ must be an embedding. Therefore, a universal globalization of $(\alpha,X)$ is a globalization, in the sense of Definition 2.19.
@@
    Let $C\mathcal{A}(\mathcal{C})$ be the full subcategory of $\mathcal{A}_{p}(\mathcal{C})$ whose objects are the categorical global actions of $\mathcal{C}$, and $I_{C}\colon C\mathcal{A}(\mathcal{C})\to\mathcal{A}_{p}(\mathcal{C})$ be the inclusion functor. Then the category $((\alpha,X)\to I_{c})$ is a full subcategory of the category $((\alpha,X)\to I)$ (recall the definition in Proposition 2.23).
@@
    From [21, Theorem 4], (2) and Proposition 2.23(i), it follows that the universal globalization $(i,\gamma,Y)$ is an initial object in the category $((\alpha,X)\to I_{C})$.
The universal globalization presented in [21] is constructed as follows. Given a partial category action $(\alpha,X)$ of $\mathcal{C}$, let $\overline{X}=\{(g,x)\colon x\in{{}_{D(g)}X}\}$. For $(g,x),(h,y)\in\overline{X}$, denote $(g,x)-(g^{\prime},x^{\prime})$ if and only if one of the following conditions is satisfied:
@@
    There is $h\in\mathcal{C}$ such that $x^{\prime}h$ is defined, $x^{\prime}h=x$, $x\in{{}_{h}X}$ and $\alpha_{h}(x)=x^{\prime}$.
@@
Let $\simeq$ be the smallest equivalence relation on $\overline{X}$ that contains the relation -- and denote $Y=X/\simeq$. For each $g\in\mathcal{C}$, let $(h,x)\in{{}_{g}\overline{X}}$ if and only if $(h,x)\simeq(h^{\prime},x^{\prime})$ for some $h^{\prime}\in\mathcal{C}$ such that $gh^{\prime}$ is defined. Denote ${{}_{g}Y}={{}_{g}\overline{X}}/\simeq$ and define $\gamma_{g}\colon{{}_{g}Y}\to Y$ by $\gamma_{g}([h,x])=[gh^{\prime},x^{\prime}]$. Then $(\gamma,Y)$ is a global action. From condition (c1), for every $x\in X$ there is $e\in\mathcal{C}_{0}$ such that $x\in{{}_{e}X}$. Define $i\colon X\to Y$ by $i(x)=[e,x]$. The triple $(i,\gamma,Y)$ is the desired universal globalization for $(\alpha,X)$.
To prove that our universal globalization $(\delta,\beta,E)$ and the one described above are isomorphic, we will use Remark 4.12(4) and the uniqueness, up to isomorphism, of initial objects.
@@
Let $(\alpha,X)$ be a categorical partial action of $\mathcal{C}$. Then the universal globalization $(\delta,\beta,E)$ of Theorem 3.7 is a categorical action of $\mathcal{C}$.
@@
We start by summarizing the construction of $(\beta,E,\delta)$ when $\mathcal{C}$ is a category and $(\alpha,X)$ is a categorical partial action of $\mathcal{C}$. The relation $R$ defined in Lemma 3.2 leads to $gRh$ if and only if $D(g)=D(h)$. Denote by $\rho_{g}=\rho_{D(g)}$ the equivalence class of $g$. Then, for each $e\in\mathcal{C}_{0}$, we have
=${{}_{\rho_{e}}}=\bigcup\left(\{{{}_{h}X}\colon D(h)=e\}\cup\{X_{h}\colon R(h)=e\}\right).$
@@
Let $E=D/\simeq$, where $\simeq$ is the equivalence relation given in Definition 3.3. For an element $(a,x)\in D$ we will denote its equivalence class by $[a,x]\in E$. Now we prove conditions (C1), (C2) and (C3).
First, let $(g,x)\in D$ with $g\in\mathcal{C}$. Since the composition $R(g)g$ is defined, we have $[g,x]\in{{}_{R(g)}E}$ by definition. Now, let $(\delta,x)\in D$. Since $(\alpha,X)$ is a categorical partial action, it follows from (C1) that $x\in X_{g}$ for some $g\in\mathcal{C}$. Since $R(g)g=g$, Lemma 4.8(ii) implies $X_{g}\subseteq X_{R(g)}$ and, from (C2), we obtain $X_{R(g)}={{}_{R(g)}X}$. Therefore $(R(g),x)\in D$ and $(\delta,x)\sim(R(g),x)$ by (1). From the first case, we have $[\delta,x]=[R(g),x]\in{{}_{R(g)}E}$. Therefore $E=\cup_{e\in\mathcal{C}_{0}}\ {{}_{e}E}$.
Let $[a,x]\in{{}_{e}E}$ for some $e\in\mathcal{C}_{0}$. Then either $[a,x]=[t,y]$ for some $t\in\mathcal{C}$ such that $R(t)=e$; or $a=\delta$ and $x\in{{}_{e}X}$. In the first case we have
@@
Therefore $\beta_{e}=id_{{}_{e}X}$. This proves (C2). Furthermore, we obtain that ${{}_{e}E}=E_{e}$ and, hence, $E=\cup_{e\in\mathcal{C}_{0}}\ {{}_{e}E}=\cup_{e\in\mathcal{C}_{0}}E_{e}$. This proves (C1). Condition (C3) follows from $(\beta,E)$ being a global action. In fact, for each $g\in\mathcal{C}$, the unique $e\in\mathcal{C}_{0}$ such that $(g,e)\in\mathcal{C}^{(2)}$ is $e=D(g)$. Therefore
@@
Let $(\alpha,X)$ be a categorical partial action of $\mathcal{C}$, $(\delta,\beta,E)$ be the universal globalization of Theorem 3.7, and $(i,\gamma,Y)$ be the universal globalization of [21, Theorem 4]. Then $(\beta,E)$ and $(\gamma,Y)$ are isomorphic in the category $\mathcal{A}(\mathcal{C})$.
@@
We know that $(\delta,\beta,E)$ is an initial object in the category $((\alpha,X)\to I)$. Since $(\alpha,X)$ is a categorical partial action, it follows from Proposition 4.13 that $(\beta,E)$ is a categorical global action. Therefore $(\delta,\beta,E)\in C\mathcal{A}(\mathcal{C})$ and, hence, $(\delta,\beta,E)$ is an initial object in the category $((\alpha,X)\to I_{C})$.

### v2

@@ 4. Application in semigroups and categories
${{}_{s}Y}=(\varphi\circ\varphi^{-1})({{}_{s}Y})\subseteq\varphi({{}_{s}X})\subseteq{{}_{s}Y}.$
That is, $\varphi({{}_{s}X})={{}_{s}Y}$. Hence,
@@
Consequently, for each $s\in\mathcal{S}$, we have ${{}_{s}E}=E\setminus\iota(X_{0})$. On the other hand, since $(i,\gamma,\mathcal{S}^{1}\otimes X)$ is a globalization of $(\alpha,X)$, there is a unique morphism of partial actions $\varphi\colon(\beta,E)\to(\gamma,\mathcal{S}^{1}\otimes X)$ such that $\varphi\circ\iota=i$. This morphism is defined by
@@
Therefore, $\varphi\colon E\to\mathcal{S}^{1}\otimes X$ is a bijection. However, the inverse function $\varphi^{-1}\colon\mathcal{S}^{1}\otimes X\to E$ may not be a morphism of partial actions, since in $(\gamma,\mathcal{S}^{1}\otimes X)$ every element acts on $i(X_{0})=\varphi(\iota(X_{0}))$ as the identity function, while in $(\beta,E)$ no element can act on $\iota(X_{0})$.
We end this section noticing that the fact that any degenerate partial semigroup action admits a non-degenerate globalization that provides an initial object in the category $((\alpha,X)\downarrow I^{\ast})$ is a convenient property of semigroups. Indeed, if $(\beta,Y)$ is a non-degenerate global action of $\mathcal{S}$, then ${{}_{s}Y}=Y$ for every $s\in\mathcal{S}$. Therefore, in order to define a morphism $\Phi\colon(\beta,E)\to(\theta,Y)$, there is no need to verify that $\Phi({{}_{s}E})\subseteq{{}_{s}Y}$. The next example illustrates, in the context of semigroupoids, how a degenerate partial action may fail to have such a non-degenerate globalization.
@@
${{}_{e}Y}$ $=\{0,1,2\},$ $\beta_{e}(0)=0,\ \beta_{e}(1)=1,\ \beta_{e}(2)=0,$
${{}_{f}Y}$ $=\{0,1,3\},$ $\beta_{f}(0)=0,\ \beta_{f}(1)=1,\ \beta_{f}(3)=0.$
@@
In this case $(\alpha,X)$ is a degenerate global action of $S$. We know that $i\colon X\to Y$, given by $i(1)=1$ and $i(2)=2$ is an embedding. Notice that $j\colon X\to Y$, given by $j(1)=1$ and $j(2)=3$, is also an embedding.
Suppose that $(\iota,\theta,E)$ is a non-degenerate globalization of $(\alpha,X)$. Since $S^{s}\neq\emptyset$ for every $s\in S$, we obtain from Lemma 2.10(ii) that $E_{s}\subseteq{{}_{t}E}$, for some $t\in S^{s}$. Therefore, $E\subseteq\cup_{s\in S}{{}_{s}E}$. Now, the embedding $\iota\colon X\to E$ must satisfy $\iota(2)\in{{}_{e}E}$ or $\iota(2)\in{{}_{f}E}$. Suppose further that $(\iota,\theta,E)$ is initial object in the category $((\alpha,X)\downarrow I^{\ast})$. Then there must be two morphisms of partial actions, $i^{\prime},j^{\prime}\colon(\theta,E)\to(\beta,Y)$, satisfying $i^{\prime}\circ\iota=i$ and $j^{\prime}\circ\iota=j$. Therefore,
$i^{\prime}(\iota(2))=i(2)=2\in{{}_{e}Y}\setminus{{}_{f}Y}\quad\text{and}\quad j^{\prime}(\iota(2))=j(2)=3\in{{}_{f}Y}\setminus{{}_{e}Y}.$
Since $i^{\prime}$ is a morphism of partial actions, if $\iota(2)\in{{}_{f}E}$ then $i^{\prime}(\iota(2))\in{{}_{f}Y}$, a contradiction. Hence, it must be $\iota(2)\in{{}_{e}E}$. But in this case, as $j^{\prime}$ is a morphism of partial actions, it must be $j^{\prime}(\iota(2))\in{{}_{e}Y}$, a contradiction. Therefore, the degenerate partial action $(\alpha,X)$ does not admit a non-degenerate globalization that provides an initial object for $((\alpha,X)\downarrow I^{\ast})$.
@@
In this subsection, we investigate the relation between partial category actions and partial semigroupoid actions of a category and show that our construction of universal globalization coincides with the one in [24, §3].
Throughout this subsection, $\mathcal{C}$ will denote a category (recall our conventions in Definition 2.3). The following characterization of partial category actions was given in [24, Proposition 8].
@@
A partial action of $\mathcal{C}$ on a set $X$ is a pair $(\{{{}_{g}X}\}_{g\in\mathcal{C}},\{\alpha_{g}\}_{g\in\mathcal{C}})$ consisting of a collection $\{{{}_{g}X}\}_{g\in\mathcal{C}}$ of subsets of $X$ and a collection $\{\alpha_{g}\}_{g\in\mathcal{C}}$ of functions $\alpha_{g}\colon{{}_{g}X}\to X$ such that:
@@
    $X=\cup_{e\in\mathcal{C}_{0}}{{}_{e}X}$ and $\alpha_{e}=id_{{}_{e}X}$, for all $e\in\mathcal{C}_{0}$;
@@
Compared to the original characterization, we omit the collection of subsets $\{X_{g}\}_{g\in\mathcal{C}}$, since it is implicitly defined as the images of the collection of functions $\{\alpha_{g}\}_{g\in\mathcal{C}}$, and interchange the roles of $X_{g}$ and $_{g}X$}. This change aims the comparison with partial group actions [11, Definition 1.2], where $\alpha_{g}\colon X_{g^{-1}}\to X_{g}$, or in our notation, $X_{g}=X_{g}$ and ${{}_{g}X}=X_{g^{-1}}$. It is straightforward from (c1) that any partial category action is non-degenerate.
Similarly to the relation between semigroups and monoids ([19, Lemma 2.4]), there is a particular class of partial semigroupoid actions that correspond to partial category actions. To describe it, we need the notion of an identity element in a semigroupoid.
@@
Notice that $S_{0}$ may be the empty set. Furthermore, for each $s\in S$, there exists at most one identity $e\in S^{s}$. Indeed, if $e,f\in S_{0}\cap S^{s}$, then $s=se=(sf)e$. Hence, $fe$ is defined, and since both $e$ and $f$ are identities, we conclude that $e=fe=f$.
@@
Let us start with (ii). Suppose that $(e,s)\in S^{(2)}$. Then $es=s$, and by (P1) we obtain $\alpha_{s}^{-1}({{}_{e}X}\cap X_{s})={{}_{s}X}\cap{{}_{es}X}={{}_{s}X}$. Applying $\alpha_{s}$ to this equality, we obtain ${{}_{e}X}\cap X_{s}=X_{s}$. Therefore $X_{s}\subseteq{{}_{e}X}$. Let $x\in{{}_{s}X}={{}_{s}X}\cap{{}_{es}X}$. Then condition (P2) implies $(\alpha_{e}\circ\alpha_{s})(x)=\alpha_{es}(x)=\alpha_{s}(x)$. That is, for every $y=\alpha_{s}(x)\in X_{s}$ we have $\alpha_{e}(y)=y\in X_{e}$. Thus, $X_{s}\subseteq{{}_{e}X}\cap X_{e}$.
@@
Let $S$ be a semigroupoid and $(\alpha,X)$ be a partial semigroupoid action. We say that $(\alpha,X)$ is a categorical partial action of $S$ if the following conditions hold:
@@
Assume (i). To prove (ii) it remains to show (C1). From the second part of (c1) we have $\alpha_{e}=id_{{}_{e}X}$, for every $e\in\mathcal{C}_{0}$. That is, ${{}_{e}X}=X_{e}$. From the first part of (c1) we obtain
@@
Suppose that (G2) holds, that is, ${{}_{h}X}={{}_{gh}X}$, whenever the composition $gh$ is defined. In particular, for $h=D(g)$, the composition $gD(g)$ exists and, hence, ${{}_{D(g)}X}={{}_{gD(g)}X}={{}_{g}X}$. This proves (c4). Assume (c4) and notice that, for $g\in\mathcal{C}$,
@@
Therefore, if $\mathcal{C}^{g}=\mathcal{C}^{h}$, then $D(g)=D(h)$. In this case ${{}_{g}X}={{}_{D(g)}X}={{}_{h}X}$. This proves that (G1) is satisfied. ∎
@@
The notion of globalization used in [24, Definition 11] is more general than ours. The term globalization means any morphism of partial actions $\iota\colon(\alpha,X)\to(\beta,Y)$ where $(\beta,Y)$ is a global action, while we require $\iota$ to be an embedding. On the other hand, universal globalization means a globalization $(\iota,\beta,Y)$ such that, for every morphism $\varphi\colon(\alpha,X)\to(\theta,Z)$ where $(\theta,Z)$ is a global action, there is a unique morphism $\Phi\colon(\beta,Y)\to(\theta,Z)$ satisfying $\Phi\circ\iota=\phi$. Apart from the definition of globalization, the idea of universal globalization agree with ours.
@@
    In [24, Theorem 4] it has been proved that $(\alpha,X)$ admits a universal globalization $(i,\gamma,Y)$. Furthermore, [24, Remark 22] shows that $i$ is, in fact, an embedding.
@@
    From (1) and Proposition 2.23(ii), it follows that, for any universal globalization $(\iota,\beta,Z)$ of $(\alpha,X)$, $\iota$ must be an embedding. Therefore, a universal globalization of $(\alpha,X)$ is a globalization, in the sense of Definition 2.19.
@@
    Let $C\mathcal{A}(\mathcal{C})$ be the full subcategory of $\mathcal{A}_{p}(\mathcal{C})$ whose objects are the categorical global actions of $\mathcal{C}$, and $I_{C}\colon C\mathcal{A}(\mathcal{C})\to\mathcal{A}_{p}(\mathcal{C})$ be the inclusion functor. Then the category $((\alpha,X)\downarrow I_{C})$ is a full subcategory of the category $((\alpha,X)\downarrow I)$ (recall the definition in Proposition 2.23).
@@
    From [24, Theorem 4], (2) and Proposition 2.23(i), it follows that the universal globalization $(i,\gamma,Y)$ is an initial object in the category $((\alpha,X)\downarrow I_{C})$.
The universal globalization from [24] is constructed as follows. Given a partial category action $(\alpha,X)$ of $\mathcal{C}$, let $\overline{X}=\{(g,x)\colon x\in{{}_{D(g)}X}\}$. For $(g,x),(g^{\prime},x^{\prime})\in\overline{X}$, denote $(g,x)-(g^{\prime},x^{\prime})$ if and only if one of the following conditions is satisfied:
@@
    There is $h\in\mathcal{C}$ such that $g^{\prime}h$ is defined, $g^{\prime}h=g$, $x\in{{}_{h}X}$ and $\alpha_{h}(x)=x^{\prime}$.
@@
Let $\simeq$ be the smallest equivalence relation on $\overline{X}$ that contains the relation - and denote $Y=X/\simeq$. For each $g\in\mathcal{C}$, let $(h,x)\in{{}_{g}\overline{X}}$ if and only if $(h,x)\simeq(h^{\prime},x^{\prime})$ for some $h^{\prime}\in\mathcal{C}$ such that $gh^{\prime}$ is defined. Denote ${{}_{g}Y}={{}_{g}\overline{X}}/\simeq$ and define $\gamma_{g}\colon{{}_{g}Y}\to Y$ by $\gamma_{g}([h,x])=[gh^{\prime},x^{\prime}]$. Then $(\gamma,Y)$ is a global action. From condition (c1), for every $x\in X$ there is $e\in\mathcal{C}_{0}$ such that $x\in{{}_{e}X}$. Define $i\colon X\to Y$ by $i(x)=[e,x]$. The triple $(i,\gamma,Y)$ is the desired universal globalization for $(\alpha,X)$.
To prove that our universal globalization $(\iota,\beta,E)$ and the one described above are isomorphic, we will use Remark 4.12(4) and the uniqueness, up to isomorphism, of initial objects.
@@
Let $(\alpha,X)$ be a categorical partial action of $\mathcal{C}$. Then the universal globalization $(\iota,\beta,E)$ of Theorem 3.7 is a categorical action of $\mathcal{C}$.
@@
We start by summarizing the construction of $(\iota,\beta,E,)$ when $\mathcal{C}$ is a category and $(\alpha,X)$ is a categorical partial action of $\mathcal{C}$. The relation $R$ defined in Lemma 3.2 leads to $gRh$ if and only if $D(g)=D(h)$. Denote by $\rho_{g}=\rho_{D(g)}$ the equivalence class of $g$. Then, for each $e\in\mathcal{C}_{0}$, we have
${{}_{\rho_{e}}X}=\bigcup\left(\{{{}_{h}X}\colon D(h)=e\}\cup\{X_{h}\colon R(h)=e\}\right).$
@@
Let $E=D/\approx$, where $\approx$ is the equivalence relation given in Definition 3.3. For an element $(a,x)\in D$ we will denote its equivalence class by $a,x]\in E$. Now we prove conditions (C1), (C2) and (C3).
First, let $(g,x)\in D$ with $g\in\mathcal{C}$. Since the composition $R(g)g$ is defined, we have $g,x]\in{{}_{R(g)}E}$ by definition. Now, let $(\delta,x)\in D$. Since $(\alpha,X)$ is a categorical partial action, it follows from (C1) that $x\in X_{g}$ for some $g\in\mathcal{C}$. Since $R(g)g=g$, Lemma 4.8(ii) implies $X_{g}\subseteq X_{R(g)}$ and, from (C2), we obtain $X_{R(g)}={{}_{R(g)}X}$. Therefore $(R(g),x)\in D$ and $(\delta,x)\sim(R(g),x)$ by (1). From the first case, we have $\delta,x]=[R(g),x]\in{{}_{R(g)}E}$. Therefore $E=\cup_{e\in\mathcal{C}_{0}}\ {{}_{e}E}$.
Let $a,x]\in{{}_{e}E}$ for some $e\in\mathcal{C}_{0}$. Then either $[a,x]=[t,y]$ for some $t\in\mathcal{C}$ such that $R(t)=e$; or $a=\delta$ and $x\in{{}_{e}X}$. In the first case we have
@@
Therefore $\beta_{e}=id_{{}_{e}X}$. This proves (C2). Furthermore, we obtain that ${{}_{e}E}=E_{e}$ and, hence, $E=\cup_{e\in\mathcal{C}_{0}}\ {{}_{e}E}=\cup_{e\in\mathcal{C}_{0}}E_{e}$. This proves (C1). Condition (C3) follows from $(\beta,E)$ being a global action. In fact, for each $g\in\mathcal{C}$, the unique $e\in\mathcal{C}_{0}$ such that $(g,e)\in\mathcal{C}^{(2)}$ is $e=D(g)$. Therefore,
@@
Let $(\alpha,X)$ be a categorical partial action of $\mathcal{C}$, $(\iota,\beta,E)$ be the universal globalization of Theorem 3.7, and $(i,\gamma,Y)$ be the universal globalization of [24, Theorem 4]. Then $(\beta,E)$ and $(\gamma,Y)$ are isomorphic in the category $\mathcal{A}(\mathcal{C})$.
@@
We know that $(\iota,\beta,E)$ is an initial object in the category $((\alpha,X)\downarrow I)$. Since $(\alpha,X)$ is a categorical partial action, it follows from Proposition 4.13 that $(\beta,E)$ is a categorical global action. Therefore $(\iota,\beta,E)\in C\mathcal{A}(\mathcal{C})$ and, hence, $(\iota,\beta,E)$ is an initial object in the category $((\alpha,X)\downarrow I_{C})$.